A bond matures on today’s date in 5 years. It has a coupon rate of 9%. The bond pays coupons annually and its yield to maturity is 11%. What is the bond's value?
Answer #1

Annual coupon=1000*9%=90

Hence bond value=Annual coupon*Present value of annuity factor(11%,5)+1000*Present value of discounting factor(11%,5)

=90*3.69589702+1000*0.593451328

=$926.08(Approx)

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=90[1-(1.11)^-5]/0.11

=90*3.69589702

2.Present value of discounting factor=1000/1.11^5

=1000*0.593451328